### v1

# Ophiuchus: Scalable Modeling of Protein Structures through Hierarchical Coarse-graining SO(3)-Equivariant Autoencoders

Abstract
Three-dimensional native states of natural proteins display recurring and hierarchical patterns. Yet, traditional graph-based modeling of protein structures is often limited to operate within a single fine-grained resolution, and lacks hourglass neural architectures to learn those high-level building blocks. We narrow this gap by introducing Ophiuchus, an SO(3)-equivariant coarse-graining model that efficiently operates on all heavy atoms of standard protein residues, while respecting their relevant symmetries. Our model departs from current approaches that employ graph modeling, instead focusing on local convolutional coarsening to model sequence-motif interactions in log-linear length complexity. We train Ophiuchus on contiguous fragments of PDB monomers, investigating its reconstruction capabilities across different compression rates. We examine the learned latent space and demonstrate its prompt usage in conformational interpolation, comparing interpolated trajectories to structure snapshots from the PDBFlex dataset. Finally, we leverage denoising diffusion probabilistic models (DDPM) to efficiently sample readily-decodable latent embeddings of diverse miniproteins. Our experiments demonstrate Ophiuchus to be a scalable basis for efficient protein modeling and generation.
1 Introduction
Proteins form the basis of all biological processes and understanding them is critical to biological discovery, medical research and drug development. Fortunately, their three-dimensional structures often display modular organization across multiple scales, making them promising candidates for modeling in motif-based design spaces [Bystroff & Baker (1998); Mackenzie & Grigoryan (2017); Swanson et al. (2022)]. Harnessing these coarser, lower-frequency building blocks is of great relevance to the investigation of the mechanisms behind protein evolution, folding and dynamics [Mackenzie et al. (2016)], and may be instrumental in enabling more efficient computation on protein structural data through coarse and latent variable modeling [Kmiecik et al. (2016); Ramaswamy et al. (2021)].
Deep learning models have been shown to produce representations that capture high-level patterns of the data [Olah et al. (2018)]. Recent developments in learning architectures to protein structures demonstrates the remarkable capabilities of neural models in the domain of protein modeling and design [Jumper et al. (2021); Baek et al. (2021b); Ingraham et al. (2022); Watson et al. (2022)]. Still, current state-of-the-art learning architectures often fail to incorporate inductive biases necessary for directly representing high-level components, such as coarsening of representations over the data domain.
To fill this gap, we introduce Ophiuchus, a deep convolutional model that captures joint encodings of sequence-structure motifs of all-atom protein structures. Our model contracts representations over the sequence to produce compressed encodings in hierarchical fashion. To capture three-dimensional geometry effectively, we model features as irreducible representations of SO(3), introducing novel roto-translational equivariant neural architectures for multi-layer hourglass-shaped autoencoder.
Figure 1: Coarsening a Three-Dimensional Sequence: (a) Ophiuchus takes atoms directly as representations of SO(3). Non-scalar signals are taken relative to the $\mathbf{C}_{\alpha}$ position of the corresponding residue. Only 3D vector $\mathbf{V}^{l=1}$ components are shown. ($\ast$) Unorderable atoms are encoded in a permutation invariant way. (b) Our proposed model uses three building blocks to transform those vectors and their anchors into coarser representations of coordinates. (c) We train a deep bottlenecking model by iteratevily coarsening and refining across the sequence axis.
We ablate the reconstruction performance of different network layouts and compression settings. We investigate the learned latent space and show it to produce rich representations that are directly employable in structural interpolation. Finally, we deploy a diffusion model on the latent space of Ophiuchus to investigate the efficacy of a coarse-to-fine method to protein structures.
Our main contributions are the following:
  • Deep Protein Autoencoding: We introduce a new autoencoding model for protein sequence-structure representation, and offer an in-depth examination of its architecture through comprehensive ablation benchmarks. We showcase the unique effectiveness and versatility of our model’s latent geometric representations in structural latent interpolation tasks.
  • Neural Architecture for Proteins: We propose novel learning algorithms for efficiently operating on all standard protein heavy atoms while preserving the natural permutation invariance of particular side-chain atoms. We introduce numerous methods for coarsening, refining and mixing of protein sequence-structure, employing irreducible representations of SO(3) to contract and spawn 3D coordinates directly from features.
  • All-Atom Protein Latent Diffusion: We explore a novel generative approach to proteins by performing latent diffusion exclusively on geometric feature representations. We train an efficient all-atom generative model for miniproteins, and provide diverse benchmarks to assess its performance.
2 Background and Related Work
2.1 Modularity and Hierarchy in Proteins
Protein sequences and structures display significant degrees of modularity. [Vallat et al. (2015)] introduces a library of common super-secondary structural motifs (Smotifs), while [Mackenzie et al. (2016)] shows protein structural space to be efficiently describable by small tertiary alphabets (TERMs). Motif-based methods have been successfully used in protein folding and design [Bystroff & Baker (1998); Li et al. (2022)].
2.2 Symmetries in Neural Architecture for Biomolecules
Learning algorithms greatly benefit from proactively exploiting symmetry structures present in their data domain [Bronstein et al. (2021); Smidt (2021)]. In this work, we investigate three relevant symmetries for the domain of protein structures:
Euclidean Equivariance of Coordinates and Feature Representations. Neural models equipped with roto-translational (Euclidean) invariance or equivariance have been shown to outperform competitors in molecular and point cloud tasks [Townshend et al. (2022); Miller et al. (2020); Deng et al. (2021)]. Similar results have been extensively reported across different structural tasks of protein modeling [Liu et al. (2022); Jing et al. (2021)]. Our proposed model takes advantage of Euclidean equivariance both in processing of coordinates and in its internal feature representations, which are composed of scalars and higher order geometric tensors [Thomas et al. (2018); Weiler et al. (2018)].
Translation Equivariance of Sequence: One-dimensional Convolutional Neural Networks (CNNs) have been demonstrated to successfully model protein sequences across a variety of tasks Karydis (2017); Hou et al. (2018); Lee et al. (2019); Yang et al. (2022). These models capture sequence-motif representations that are equivariant to translation of the sequence. However, sequential convolution is less common in architectures for protein structures, which are often cast as Graph Neural Networks (GNNs) [Zhang et al. (2021)]. Notably, [Fan et al. (2022)] proposes a CNN network to model the regularity of one-dimensional sequences along with three-dimensional structures, but they restrict their layout to coarsening. In this work, we further integrate geometry into sequence by directly using three-dimensional vector feature representations and transformations in 1D convolutions. We use this CNN to investigate an unsupervised autoencoding approach to protein structures.
Permutation Invariances of Side-chain Atoms. In order to capture the permutable ordering of atoms, neural models of molecules are often implemented with permutation-invariant GNNs [Wieder et al. (2020)]. Nevertheless, protein structures are sequentially ordered, and most standard side-chain atoms are readily orderable, with four exceptions of residues [Jumper et al. (2021)]. We use this fact to design an efficient approach to directly model all heavy atoms of proteins, proposing a method to parse atomic positions and symmetries through roto-translational equivariant features.
2.3 Unsupervised Learning of Proteins
Unsupervised techniques for capturing protein sequence and structure have witnessed remarkable advancements in recent years [Lin et al. (2023); Elnaggar et al. (2023); Zhang et al. (2022)]. Amongst unsupervised methods, autoencoder models learn to produce efficient low-dimensional representations using an informational bottleneck. These models have received significant attention for enabling learning of coarse representations of molecules [Wang & Gómez-Bombarelli (2019); Yang & Gómez-Bombarelli (2023); Winter et al. (2021); Wehmeyer & Noé (2018); Ramaswamy et al. (2021)], and have been successfully deployed to diverse protein tasks of modeling and sampling [Wang et al. (2022); Eguchi et al. (2020); Lin et al. (2021); Mansoor et al. (2023)].
2.4 Denoising Diffusion for Proteins
Denoising Diffusion Probabilistic Models (DDPMs) [Sohl-Dickstein et al. (2015); Ho et al. (2020)] have found widespread adoption through diverse architectures for generative sampling of protein structures. Chroma [Ingraham et al. (2022)] trains random graph message passing through roto-translational invariant features, while RFDiffusion [Watson et al. (2022)] fine-tunes pretrained folding model RoseTTAFold [Baek et al. (2021a)] to denoising, employing SE(3)-equivariant transformers in structural decoding [Fuchs et al. (2020)]. [Yim et al. (2023); Anand & Achim (2022)] generalize denoising diffusion to frames of reference, employing Invariant Point Attention [Jumper et al. (2021)] to model three dimensions. More recently, [Fu et al. (2023)] designs a latent diffusion model by employing EGNNs [Satorras et al. (2022)] to contract the coordinate dimension into a smaller space. In contrast, our model uses SO(3)-equivariant feature representations to produce increasingly richer structural features from autoencoded coordinates. We perform generation directly on these features to demonstrate a latent diffusion generative model.
3 The Ophiuchus Architecture
We treat a protein as an ordered list $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})_{1:N}$ of $N$ nodes, each with an anchoring position $\mathbf{P}_{i}\in\mathbb{R}^{1\times 3}$ and a collection of irreducible representations of SO(3): $\mathbf{V}_{i}^{l}\in\mathbb{R}^{d\times(2l+1)}$ for integers $l\in[0,{l_{\max}}]$, $i\in[1,N]$. These representations are directly produced from sequence labels and all-atom positions, as we describe in the Atom Encoder/Decoder sections. Our model then acts with diverse operations on these representations. Self-Interaction updates vector representations independently, mixing within same-anchored vectors of different orders. In order to mix across the domain, Sequence Convolution updates $K$-sized windows of the list simultaneously, message-passing within sequence neighbors. Finally, Spatial Convolutions employ permutation invariant geometric messages to update the representations and coordinates of spatial neighbors. These modules are composed to build a coarsening-refining encoder-decoder hourglass architecture.
3.1 All-Atom Atom Encoder and Decoder
The input format of our modeling starts at the residue level. Given a particular $i$-th residue, let $\mathbf{R}_{i}\in\mathcal{R}$ denote its residue label, $\mathbf{P}_{i}^{\alpha}\in\mathbb{R}^{1\times 3}$ denote the global position of its alpha carbon ($\mathbf{C}_{\alpha}$), and $\mathbf{P}^{\ast}_{i}$ the position of all other heavy atoms relative to $\mathbf{C}_{\alpha}$. We produce initial residue representations $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})^{0}_{i}$ by setting anchors $\mathbf{P}_{i}^{0}=\mathbf{P}^{\alpha}_{i}$, scalar representations $\mathbf{P}_{i}^{0,l=0}=\textrm{Embed}(\mathbf{R}_{i})$ , and vector representations $\mathbf{V}_{i}^{0,l>0}$ to explicitly encode relative atomic positions $\mathbf{P}^{\ast}_{i}$. In particular, provided the residue label $\mathbf{R}_{i}$, heavy atoms of most standard protein residues are readily orderable and may be directly processed as a stack of $\mathbf{V}^{l=1}$ relative vector positions.
However, some of the standard residues do present two-permutations within their atom configurations (i.e. they are not orderable). To handle these cases, we map the set of two relative vectors into a pair $(\mathbf{V}^{l=1},\mathbf{V}^{l=2})$ where $\mathbf{V}^{l=1}$ is the center and $\mathbf{V}^{l=2}$ is the unsigned difference. This signal is invariant to corresponding atomic two-flips, while still carrying relevant information on positioning and angularity. Decoding is handled similarly. Refer to Appendix A for further details.
Figure 2: Building Blocks of Ophiuchus: (a) Atom Encoder and (b) Atom Decoder enable the model to directly take and produce atomic coordinates. (c) Self Interaction updates representations internally, across different vector orders $l$. (d) Sequence Convolution and (e) Transpose Sequence Convolution communicate sequence neighbors and produce coarser and finer representations, respectively. (d) Spatial Convolution interacts spatial neighbors. (g) Hourglass model: we compose those modules to build encoder and decoder models, stacking them into an autoencoder.
Input: $\mathbf{C}_{\alpha}$ Position $\mathbf{P}^{\alpha}\in\mathbb{R}^{1\times 3}$
Input: Heavy Atom Relative Positions $\mathbf{P}^{\ast}\in\mathbb{R}^{n\times 3}$
Input: Residue Label $\mathbf{R}\in\mathcal{R}$
Output: Latent Representation $(\mathbf{P},\mathbf{V}^{l=0:2})$
$\mathbf{P}\leftarrow\mathbf{P}^{\alpha}$ $\mathbf{V}^{l=0}\leftarrow\textrm{Embed}(\mathbf{R})$ $\mathbf{V}^{l=1}\leftarrow\textrm{UnorderableCenters}(\mathbf{R},\mathbf{P}^{\ast})$ $\mathbf{V}^{l=2}\leftarrow\textrm{UnorderableDifferences}(\mathbf{R},\mathbf{P}^{\ast})$ $\mathbf{V}^{l=1}\leftarrow\mathbf{V}^{l=1}\oplus\textrm{Orderable}(\mathbf{R},\mathbf{P}^{\ast})$ return $(\mathbf{P},\mathbf{V}^{l=0:2})$
Algorithm 1 All-Atom Encoding
Input: Latent Representation $(\mathbf{P},\mathbf{V}^{l=0:2})$
Output: $\mathbf{C}_{\alpha}$ Position $\mathbf{P}^{\alpha}\in\mathbb{R}^{1\times 3}$
Output: Heavy Atom Relative Positions $\mathbf{P}^{\ast}\in\mathbb{R}^{n\times 3}$
Output: Residue Label $\mathbf{R}\in\mathcal{R}$
$\mathbf{P}^{\alpha}\leftarrow\mathbf{P}$ $\mathbf{R}\leftarrow\mathrm{Softmax}(\mathbf{V}^{l=0})$ $\mathbf{P}^{\ast}\leftarrow\mathrm{Orderable}(\mathbf{R},\mathbf{V}^{l=1})$ $\mathbf{P}^{\ast}\leftarrow\mathbf{P}^{\ast}\oplus\mathrm{Unorderable}(\mathbf{R},\mathbf{V}^{l=1},\mathbf{V}^{l=2})$ return $(\mathbf{P}^{\alpha},\mathbf{P}^{\ast},\mathbf{R})$
Algorithm 2 All-Atom Decoding
This processing makes Ophiuchus strictly blind to ordering flips within permutable atoms, while still enabling it to operate directly on all heavy atoms in an efficient, stacked representation.
3.2 Self-Interaction
Our Self-Interaction transform draws inspiration from MACE [Batatia et al. (2022)]. This transformation updates the feature vectors $\mathbf{V}^{0:{l_{\max}}}_{i}$ localized at the same anchoring node $i$. Importantly, it blends feature vectors $\mathbf{V}^{l}$ of varying orders $l$ by employing tensor products of the features with themselves. We offer two implementations of these tensor products to cater to different computational needs. For a comprehensive explanation, please refer to Appendix B.
Input: Latent Representation $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
$\tilde{\mathbf{V}}^{0:{l_{\max}}}\leftarrow\mathbf{V}^{0:{l_{\max}}}$  normal-$\triangleright$ Save For Later
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\mathbf{V}^{0:{l_{\max}}}\oplus\left(\mathbf{V}^{0:{l_{\max}}}\right)^{\otimes 2}$  normal-$\triangleright$ Tensor Product with itself
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\mathrm{Linear}(\mathbf{V}^{0:{l_{\max}}})$ $\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{MLP}(\mathbf{V}^{l=0})\cdot\mathbf{V}^{0:{l_{\max}}}$  normal-$\triangleright$ Scalar-Irrep Multiplication
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\mathbf{V}^{0:{l_{\max}}}+\tilde{\mathbf{V}}^{0:{l_{\max}}}$  normal-$\triangleright$ If the Multiplicities Match
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Norm}(\mathbf{V}^{0:{l_{\max}}})$  normal-$\triangleright$ Normalize $l$-Wise
return $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
Algorithm 3 Self-Interaction
3.3 Sequence Convolution
We introduce a one-dimensional roto-translational equivariant convolutional kernel designed to operate on representations of SO(3) and $\mathbb{R}^{3}$. Given a kernel window size $K$ and stride $S$, we concatenate representations $\mathbf{V}^{0:{l_{\max}}}_{i-\frac{K}{2}:i+\frac{K}{2}}$ with the same $l$ value. Additionally, we include normalized relative vectors between anchoring positions $\mathbf{P}_{i-\frac{K}{2}:i+\frac{K}{2}}$. Following conventional CNN architectures, this concatenated representation undergoes a linear transformation. The scalars in the resulting representation are then converted into weights, which are used to combine window coordinates into a new coordinate. To ensure translational equivariance, these weights are constrained to sum to one.
Input: Window of Latent Representations $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})_{1:K}$
$w_{1:K}\leftarrow\textrm{Softmax}\Big{(}\textrm{MLP}\big{(}\mathbf{V}^{0}_{1:K}\big{)}\Big{)}$   normal-$\triangleright$ Such that $\sum_{k}w_{k}=1$
$\mathbf{P}\leftarrow\sum_{k=1}^{K}w_{k}\mathbf{P}_{k}$ $\tilde{\mathbf{P}}_{1:K,1:K}\leftarrow\bigoplus_{i,j}\frac{\mathbf{P}_{i}-\mathbf{P}_{j}}{\|\mathbf{P}_{i}-\mathbf{P}_{j}\|}$   normal-$\triangleright$ All relative vectors
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Linear}\Big{(}\mathbf{V}^{0:{l_{\max}}}_{1:K}\oplus\tilde{\mathbf{P}}_{1:K,1:K}\Big{)}$ return $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
Algorithm 4 Sequence Convolution
When $S>1$, sequence convolutions reduce the dimensionality along the sequence axis, yielding mixed representations and coarse coordinates. To reverse this process, we introduce a transpose convolution algorithm that geometrically reconstructs both the sequence and its positional coordinates. For further details, please refer to Appendix C.
3.4 Spatial Convolution
Spatial convolutions update representations through message passing within k-nearest spatial neighbors. Message representations incorporate SO(3) signals from the vector difference between neighbor coordinates, and we aggregate messages with permutation-invariant means. After aggregation, we linearly transform the vector representations into a single vector that is used to update the coordinates.
Input: Latent Representations $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})_{1:N}$
Input: Output Node Index $i$
$\hat{\mathbf{V}}^{0:{l_{\max}}}\leftarrow\mathbf{V}^{0:{l_{\max}}}_{i}$  normal-$\triangleright$ Save for Later
$(\tilde{\mathbf{P}},\tilde{\mathbf{V}}^{0:{l_{\max}}})_{1:k}\leftarrow\textrm{$k$-Nearest-Neighbors}(\mathbf{P}_{i},\mathbf{P}_{1:N})$ $(\phi,R,u)_{1:k}\leftarrow\textrm{Embeddings}(\tilde{\mathbf{X}}_{1:k}-\mathbf{P}_{i})$  normal-$\triangleright$ Angular/Radial Embeddings
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Linear}\left(\left(\sum_{k}u_{k}\textrm{MLP}(R_{k})\tilde{\mathbf{V}}^{0:{l_{\max}}}_{k}\otimes\phi_{k}\right)/\left(\sum_{k}u_{k}\right)\right)$ $\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Norm}\left(\mathbf{V}^{0:{l_{\max}}}\right)$  normal-$\triangleright$ Normalize $l$-Wise
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Linear}\left(\mathbf{V}^{0:{l_{\max}}}\oplus\hat{\mathbf{V}}^{0:{l_{\max}}}\right)$ $\mathbf{P}\leftarrow\mathbf{P}_{i}+10^{-3}\;\textrm{Linear}\left(\mathbf{V}^{l=1}\right)$ return $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
Algorithm 5 Spatial Convolution
3.5 Deep Coarsening Autoencoder
We compose Space Convolution, Self-Interaction and Sequence Convolution modules to define a coarsening/refining block. The block mixes representations across all relevant axes of the domain while producing coarsened, downsampled positions and mixed embeddings. The reverse result – finer positions and decoupled embeddings – is achieved by changing the standard Sequence Convolution to its transpose counterpart. When employing sequence convolutions of stride $S>1$, we increase the dimensionality of the feature representation according to a rescaling factor hyperparameter (Table 1). We stack $L$ coarsening blocks to build a deep neural encoder $\mathcal{E}$, and symetrically $L$ refining blocks to build a decoder $\mathcal{D}$.
3.6 Reconstruction Losses
We use a number of reconstruction losses to ensure good quality of produced proteins.
Vector Map Loss. We train the model by directly comparing internal three-dimensional vector difference maps. Let $V(\mathbf{P})$ denote the internal vector map between all atoms $\mathbf{P}$ in our data, that is, $V(\mathbf{P})^{i,j}=(\mathbf{P}^{i}-\mathbf{P}^{j})\in\mathbb{R}^{N\times N\times 3}$. We then directly minimize $\textrm{HuberLoss}(V(\mathbf{P}),V(\hat{\mathbf{P}}))$ [Huber (1992)].
Cross Entropy Loss. We train the model to predict logits over alphabet $\mathcal{R}$ per residue. We then minimize the cross entropy between predicted logits and ground labels.
Chemical Losses. We incorporate $L_{2}$-norm losses for comparing bonds, angles and dihedrals between prediction and ground truth. For nonbonded atoms, a clash loss is evaluated using standard Van der Waals atomic radii.
An additional stage is employed for processing permutation symmetry breaks. Please refer to Appendix D for further details.
3.7 Implementation
We train all models on single-GPU A6000 and GTX-1080 Ti machines. We implement Ophiuchus in Jax using the python libraries e3nn-jax [Geiger & Smidt (2022)] and Haiku [Hennigan et al. (2020)].
4 Methods and Results
4.1 Architecture Ablation
To investigate the effects of different architecture layouts and coarsening rates, we train different instantiations of Ophiuchus to coarsen representations of contiguous 160-sized protein fragments from the Protein Databank (PDB) [Berman et al. (2000)]. We filter out entries tagged with resolution higher than 2.5 Å, and total sequence lengths larger than 512. We also ensure proteins in the dataset have same chirality. For ablations, the sequence convolution uses kernel size of 5 and stride 3, channel rescale factor per layer is one of {1.5, 1.7, 2.0}, and the number of downsampling layers ranges in 3-5. The initial residue representation uses 16 channels, where each channel is composed of one scalar ($l=0$) and one 3D vector ($l=1$). All experiments were repeated 3 times with different initialization seeds and data shuffles.
Table 1: Autoencoder Recovery from Compressed Representations
[…]
4.2 Latent Conformational Interpolation
Figure 3: Latent Conformational Interpolation. (a) Comparison of chemical validity metrics across interpolated structures of Nuclear Factor of Activated T-Cell (NFAT) Protein (PDB ID: 1S9K and PDB ID: 2O93). We plot results for interpolations on the original data space, on the latent space, and on the autoencoder-reconstructed data space. We note that near interpolation endpoints, latent and AE-reconstructed interpolations perform worse than direct interpolations. This is expected and is a result of the low resolution decoding of the AE. However, only for the latent interpolation the chemical deviation stays consistent, demonstrating that our latent space compactly and smoothly represents the protein space. (b) Linear interpolation of coordinates between two conformational endpoints. (c) Decoded linear latent interpolation between the same encoded states as in (b).
To demonstrate the power of Ophiuchus’s geometric latent space, we show smooth interpolation between two states of a protein structure without explicit latent regularization [Ramaswamy et al. (2021)]. We utilize PDBFlex dataset [Hrabe et al. (2016)] to pick pairs of flexible conformational snapshots of a protein. We utilize these pairs as endpoints for our interpolation. We train a large Ophiuchus reconstruction model on general PDB data. The model coarsens up to 485-residue proteins into a single high-order geometric representation using 6 convolutional downsampling layers each with kernel size 5 and stride 3. The endpoint structures are compressed into single high-order geometric representation to enable direct latent interpolation. We then compare the results of linear interpolation in this latent space against linear interpolation in the coordinate domain (Fig. 3). To determine chemical validity of intermediate states, we scan protein data to quantify average bond lengths and inter-bond angles. We then calculate $L_{2}$ deviation of bonds and angles of interpolated structures. Additionally, we measure atomic clashes by counting collisions of van der Waals radii of non-bonded atoms. Our experiment shows that the decoded structures from latent interpolation maintain a consistent profile of chemical validity, while direct interpolation in the coordinate domain disrupts it significantly. We provide an additional example in Appendix G.
[…]
4.3 Latent Diffusion
We investigate the performance of a generative model trained on the Ophiuchus latent space for sampling decodable protein structure representations. We leverage the miniprotein dataset produced by [Cao et al. (2022)] for protein binder design. The dataset consists of 66k all-atom structures of sequence length between 50 and 65, and composes of diverse folds and sequences across 5 secondary structure topologies. The small sequence sizes facilitate precise reconstruction from unified bottlenecks, enabling a generation method exclusively on latent feature representations to implicitly generate all-atom coordinates.
To that end, we employ DDPMs [Ho et al. (2020); Sohl-Dickstein et al. (2015)] on signals over SO(3) to produce decodable embeddings. First, we train deep autoencoder models for backbone and all-atom structures over the miniprotein dataset. In order to ensure that bottleneck representations $\mathbf{Z}_{0}$ are well-behaved targets for generation, we further regularize the latent space. We build a denoising network by stacking $L_{D}=7$ layers of the Ophiuchus Self-Interaction Fig. (4.a). We parameterize the diffusion model following [Salimans & Ho (2022)], and train it to recover $\mathbf{Z}_{0}$ from $\mathbf{Z}_{t}$: $\mathcal{L}_{\textrm{diffusion}}=\mathbb{E}_{{\epsilon},t}\big{[}w(\lambda_{t})||\hat{\mathbf{Z}}_{0}(\mathbf{Z}_{t})-\mathbf{Z}_{0}||^{2}_{2}\big{]}$. (Appendix H)
In Fig.(5.b) we show the distribution of TM-scores for joint sampling of sequence and all-atom structure by the diffusion model. We produce marginal distributions of generated samples Fig.(5.e) and find them to successfully approximate the densities of ground truth data. To test the robustness of joint sampling of structure and sequence, we compute self-consistency TM-Scores[Trippe et al. (2023)]. 54% of our sampled backbones have scTM scores > 0.5 compared to 77.6% of samples from RFDiffusion. We also sample a protein between 50-65 amino acids in an average 0.21s compared to 11s taken by RFDiffusion on a RTX A6000. Details can be found in Appendix H.
Figure 5: Self-Consistency Template Matching Scores for Ophiuchus Diffusion. (a) Distribution of scTM scores for 500 sampled backbone. (b) Distribution of TM-Scores of jointly sampled backbones and sequences from an all-atom diffusion model and corresponding OmegaFold models.(c) Distribution of average plDDT scores for 500 sampled backbones (d) TM-Score between sampled backbone(green) and OmegaFold structure(blue) (e) TM-Score between sampled backbone(green) and the most confident OmegaFold structure of the sequence recovered from ProteinMPNN (blue)
5 Conclusion and Future Work
In this work, we introduced a new unsupervised autoencoder model for protein structure and sequence representation. Through extensive ablation on its architecture, we quantified the trade-offs between model complexity and representation quality. We demonstrated the power of our learned representations in latent interpolation, and investigated its usage as basis for efficient latent generation of all-atom proteins. Our studies suggest Ophiuchus to provide a strong foundation for constructing state-of-the-art protein neural architectures. In future work, we will investigate scaling Ophiuchus representations and generation to larger proteins and additional molecular domains.
Appendix
Appendix A All-Atom Decoding
Given a latent representation at the residue-level, $(\mathbf{X},\mathbf{V}^{0:{l_{\max}}})_{i}$, we take $\mathbf{X}_{i}$ directly as $\hat{\mathbf{X}}_{i}^{\alpha}$ to obtain the $\mathbf{C}_{\alpha}$ position, and transform hidden scalar representations $\mathbf{V}_{i}^{l=0}$ into categorical logits over $\mathcal{R}$ to decode $\hat{\mathbf{R}}_{i}$. We then transform $\mathbf{V}_{i}^{l>0}$ to obtain the relative position vectors $\hat{\mathbf{X}}^{\ast}_{i}$. The orderable atoms are obtained from $\mathbf{V}^{l=1}$ while the unorderable pairs are obtained from $(\mathbf{V}^{l=1},\mathbf{V}^{l=2})$. To generate an unorderable pair, we treat $\mathbf{V}^{l=1}$ as the center position of the symmetric atoms – that is, a "mean vector". We then compute the square root of $\mathbf{V}^{l=2}$ to establish the relative difference between the two vectors in the unorderable pair. This method proves effective because the sign of the square root is inherently ambiguous. Both positive and negative square root values yield the same $\mathbf{V}^{l=2}$, aligning perfectly with our requirement for the vectors to be unorderable.
Appendix B Details on Self-Interaction
The objective of our Self-Interaction module is to function exclusively based on $\mathbf{V}^{0:{l_{\max}}}$, while concurrently mixing irreducible representations across various $l$ values. To accomplish this, we calculate the tensor product of the representation with itself; this operation is termed the "tensor square" and is symbolized by $\left(\mathbf{V}^{0:{l_{\max}}}\right)^{\otimes 2}$. As the channel dimensions expand, the computational complexity tied to the tensor square increases quadratically. In the case of latent diffusion models, we employ the tensor square in its original, unmodified form, accepting the computational cost as a trade-off for maintaining full expressivity. Contrastingly, for autoencoder models, we reduce the computational load by dividing the representation into smaller, independent segments along the channel dimension prior to executing the tensor square. The segmented results are subsequently concatenated to generate the updated representation $\mathbf{V}^{0:{l_{\max}}}$.
Appendix C Transpose Sequence Convolution
Given a single, coarse anchor and representation $\{\mathbf{P},\mathbf{V}^{l}\}$, we first map $\{V^{l}\}$ into $K$ new representations linearly, reshaping it appropriately. We then similarly map $\mathbf{V}^{l=1}$ linearly into relative positioning vectors $[\Delta\mathbf{P}]^{K}$, anchoring them in the previous coordinate to obtain $K$ new coordinates, $[\mathbf{P}]^{K}\leftarrow\mathbf{P}+[\Delta\mathbf{P}]^{K}$:
Input: Kernel Size $K$ and Stride $S$.
Input: The Perceptive Field of Varying Size $P$ ($\in\{1,2\}$ for $K=3$ and $S=2$)
Input: Latent Representations $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})_{1:P}$
Input: Relative Indices $\Delta_{1:P}\in[-\frac{K}{2},\frac{K}{2}]$
$\mathbf{V}^{0:{l_{\max}}}_{1:P}\leftarrow\textrm{Norm}\left(\mathbf{V}^{0:{l_{\max}}}_{1:P}\right)$  normal-$\triangleright$ Normalize $l$-Wise
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\frac{1}{P}\sum_{p=1}^{P}\textrm{Linear}\left(\Delta_{p},\mathbf{V}^{0:{l_{\max}}}_{p}\right)$  normal-$\triangleright$ Different Parameters For Each $\Delta$
$\mathbf{P}\leftarrow\frac{1}{P}\sum_{p=1}^{P}(\mathbf{P}_{p}+\textrm{Linear}\left(\Delta_{p},\mathbf{V}^{l=1}_{p}\right))$ return $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
Algorithm 6 Transpose Sequence Convolution
Windows generated through Transpose Convolutions predict anchoring positions and representations that intersect when mapped back to sequence. We resolve those superpositions by taking the mean anchor and mean representations across intersections.
Appendix D Loss Details
Our loss function minimizes differences of internal vector maps $||V(\mathbf{P})-V(\hat{\mathbf{P}})||^{1}_{1}$ between predicted and ground positions, where $V(\mathbf{P})^{i,j}=(\mathbf{P}^{i}-\mathbf{P}^{j})\in\mathbb{R}^{N\times N\times 3}$. Our output algorithm for decoding atoms produces arbitrary symmetry breaks (Appendix A) for permutation invariances of particular side-chain atoms. Because of that, the vector map loss is not directly applicable to the output of our model. Instead, we first compare vector maps internal to individual residues. We consider the alternative permutation of ground truth positions, and pick the solution which minimizes the vector map loss. We then flip ground truth for the rest to apply the loss across residues in all-atom settings. A similar approach to tackling side-chain symmetries is described by [Jumper et al. (2021)].
Appendix E Visualization of Latent Space
To visualize the learned latent space of Ophiuchus, we forward 50k samples from the training set through a large 485-length model (Figure 6). We collect the scalar component of bottleneck representations, and use t-SNE to produce 2D points for coordinates. We similarly produce 3D points and use those for coloring. The result is visualized in Figure 6. Visual inspection of neighboring points reveals unsupervised clustering of similar folds and sequences.
Figure 6: Latent Space Analysis: we qualitatively examine the unsupervised clustering capabilities of our representations through t-SNE.
Appendix F More on Ablation
In addition to all-atom ablation found in Table 1 we conducted a similar ablation study on models trained on backbone only atoms as shown in Table 2. We found that backbone only models performed slightly better on backbone reconstruction. Furthermore, in Fig. 7.a we compare the relative vector loss between ground truth coordinates and the coordinates reconstructed from the autoencoder with respect to different downsampling factors. We average over different trials and channel rescale factors. As expected, we find that for lower downsampling factors the structure reconstruction accuracy is better. In Fig 7.b we similarly plot the residue recovery accuracy with respect to different downsampling factors. Again, we find the expected result, the residue recovery is better for lower downsampling factors. Notably, the relative change in structure recovery accuracy with respect to different downsampling factors is much lower compared to the relative change in residue recovery accuracy for different downsampling factors. This suggest our model was able to learn much more efficiently a compressed prior for structure as compared to sequence, which coincides with the common knowledge that sequence has high redundancy in biological proteins. In Fig. 7.c we compare the structure reconstrution accuracy across different channel rescaling factors. Interestingly we find that for larger rescaling factors the structure reconstruction accuracy is slightly lower. However, since we trained only for 10 epochs, it is likely that due to the larger number of model parameters when employing a larger rescaling factor it would take somewhat longer to achieve similar results. Finally, in Fig. 7.d we compare the residue recovery accuracy across different rescaling factors. We see that for higher rescaling factors we get a higher residue recovery rate. This suggests that sequence recovery is highly dependant on the number of model parameters and is not easily capturable by efficient structural models.
Figure 7: Ablation Training Curves. We plot metrics across 10 training epochs for our ablated models from Table 1. (a-b) compares models across downsampling factors and highlights the tradeoff between downsampling and reconstruction. (c-d) compares different rescaling factors for fixed downsampling at 160.
Table 2: Recovery rates from bottleneck representations - Backbone only
[…]
17 [16, 24, 36] 0.34 0.81 $\pm$ 0.31 96 $\pm$ 3 81 $\pm$ 5
17 [16, 27, 45] 0.38 0.99 $\pm$ 0.45 95 $\pm$ 3 81 $\pm$ 6
17 [16, 32, 64] 0.49 1.03 $\pm$ 0.42 92 $\pm$ 4 74 $\pm$ 6
53 [16, 24, 36, 54] 0.49 0.99 $\pm$ 0.38 92 $\pm$ 5 74 $\pm$ 8
53 [16, 27, 45, 76] 0.67 1.08 $\pm$ 0.40 91 $\pm$ 6 71 $\pm$ 8
53 [16, 32, 64, 128] 1.26 1.02 $\pm$ 0.64 92 $\pm$ 9 75 $\pm$ 11
160 [16, 24, 36, 54, 81] 0.77 1.33 $\pm$ 0.42 84 $\pm$ 7 63 $\pm$ 8
160 [16, 27, 45, 76, 129] 1.34 1.11 $\pm$ 0.29 89 $\pm$ 4 69 $\pm$ 7
160 [16, 32, 64, 128, 256] 3.77 0.90 $\pm$ 0.44 94 $\pm$ 7 77 $\pm$ 9
Appendix G Additional Conformational Interpolation
We provide an additional example from PDBFlex in Fig. 8.
Figure 8: (a) Comparison of chemical Comparison of chemical validity metrics across interpolated structures of Maltose-binding periplasmic protein (PDB ID: 4JKM and PDB 6LF3). We plot results for interpolations on the original data space, on the latent space, and on the autoencoder-reconstructed data space. (b) Direct linear interpolation of coordinates. (c) Decoded linear latent interpolation between the same encoded states as in (b).
Appendix H Diffusion Details
To train an autoencoder for latent diffusion over the miniprotein library, we instantiate an Ophiuchus model with stride $S=3$ and kernel size $K=5$. The model progressively coarsens sequences across lengths [63, 31, 15, 7, 3, 1] while producing representations of dimensionalities [12, 18, 27, 40, 60, 90].
We regularize the learned latent space so that representations are amenable to the range scales of diffusion. Let $\mathbf{V}^{l}_{i}$ denote the $i$-th channel of a vector representation $\mathbf{V}^{l}$. We regularize the autoencoder latent space by optimizing radial and angular components of our vectors:
$\mathcal{L}_{\textrm{reg}}=\sum_{i}\textrm{ReLU}(1-||\mathbf{V}_{i}^{l}||^{1}_{1})+\sum_{i}\sum_{i\neq j}(\mathbf{V}_{i}^{l}\cdot\mathbf{V}_{j}^{l})$
The first term penalizes vector magnitudes larger than one, and the second term induces vectors to spread angularly. We find these regularizations to significantly help training of the denoising diffusion model. The diffusion model uses Self-Interactions with full tensor product [B].
To compute the scTM scores, we recover 8 sequences using ProteinMPNN for all sampled backbones from our diffusion model. We used a sampling temperature of 0.1 for ProteinMPNN. Unlike the original work, where the sequences where folded using AlphaFold2, we use OmegaFold[Wu et al. (2022)] similar to [Lin & AlQuraishi (2023)].The predicted structures are aligned to the original sampled backbones and TM-Score is calculated for each alignment. For the all-atom generation we calculated the TM-scores by folding the generated sequence using OmegaFold (Fig. 5.b.)